Find the x intercepts and the vertex of the parabola. (If an answer does not exist, enter DNE.)
f(x) = x2 − 16x + 48
x= (smaller value)
x= (larger value)

Answer by stanbon(75887)

Find the x intercepts and the vertex of the parabola. (If an answer does not exist, enter DNE.)
f(x) = x^2 − 16x + 48
To find the x-intercepts, solve the following::
x^2 - 16x + 48 = 0
(x-4)(x-12) = 0
x = 4 or x = 12
------
The x-value of the vertex is midway between the x-intercepts::
That x-valsue is 4+ (12-4)/2 = 4+4 = 8
f(8) = 64 - 128 + 48
f(8) = -16
----
Vertex is at (8,-16)
---------------
Cheers,
Stan H.
-------------------